OpenStudy (anonymous):

what is the sum of the geometric sequence -4, 24, -144... if there are 7 terms?

OpenStudy (primeralph):

@Jhannybean Please do this.

OpenStudy (ivettef365):

those are multiple of -6

OpenStudy (jhannybean):

formula for finding the sum of a geometric sequence is \[\large a\left ( \frac{1-r^{n}}{1-r} \right )\]\[\large {n= 7 \\r = -6\\ a = -4 }\] your "a" is the first term present, your ratio is found by dividing the preceding term by previous term, and your "n" is the number of terms you want to find in your sequence. \[r = \frac{24}{-4} = -6 \ , \ \frac{-144}{24} = -6 \, \ ...\]plugging these values in. \[\large -4 \left ( \frac{1-(-6)^{7}}{1-(-6)}\right )\]\[\large -4 \left ( \frac{279937}{7} \right )=-159964\]

OpenStudy (anonymous):

thank you very much!

OpenStudy (jhannybean):

Do you understand how it is found? it is a lot different than finding sums of geometric series.

OpenStudy (jhannybean):

Not sure if you've learned those yet...